2p(a)+p(a-1) if p(x)=x^2+3
we know what the answer is but cannot figure out the steps to get to it
helping son with algebra II homework

Answer by swincher4391(1107)

p(a) = a^2 + 3
p(a-1) = (a-1)^2 + 3 = (a^2 -2a + 1) +3 by foil.
so continuing: p(a-1) = a^2 -2a +4
Now substitute.
2*(a^2+3) + a^2 -2a +4
2a^2 + 6 +a^2 - 2a + 4 by distributive property
Now combine like terms
(2a^2 + a^2) - 2a + (6+4)
3a^2 - 2a + 10
If you have any questions about the steps presented here, email me at my private email: OnlyYourMathTutor@yahoo.com
Otherwise hope this was clear!
-Devin